An unknown compound contains the elements carbon, hydrogen and oxygen. When a 6.85g sample of the compound was analysed by combustion in pure oxygen, 11.93g of CO2 and 3.05g of H2O were obtained. Calculate the empirical formula of the compound.
1
Expert's answer
2020-05-07T14:08:29-0400

Solution:

CxHyOz - an unknown compound.

The Molar mass of CO2 is 44.01 g/mol.

The Molar mass of H2O is 18.015 g/mol.

Let's begin by finding the moles of CO2 and H2O:

n(CO2) = (11.93 g CO2) × (1 mol CO2 / 44.01 g) = 0.2711 mol CO2

n(H2O) = (3.05 g H2O) × ( 1 mol H2O / 18.015 g) = 0.1693 mol H2O


According to the law of conservation of mass:

1) CO2 → C

Therefore,

n(C) = n(CO2) = 0.2711 mol

n(C) = 0.2711 mol

m(C) = n(C) × M(C) = (0.2711 mol) × (12.01 g/mol) = 3.256 g


2) H2O → 2H

Therefore,

n(H) = 2 × n(H2O) = 2 × (0.1693 mol) = 0.3386 mol

n(H) = 0.3386 mol

m(H) = n(H) × M(H) = (0.3386 mol) × (1.008 g/mol) = 0.3413 g


3) m(CxHyOz) = m(C) + m(H) + m(O)

m(O) = m(CxHyOz) - m(C) - m(H)

m(O) = 6.85 g - 3.256 g - 0.3413 g = 3.2527 g

n(O) = m(O) / M(O) = (3.2527 g) / (15.999 g/mol) = 0.2033 mol

n(O) = 0.2033 mol


n(C) : n(H) : n(O) = 0.2711 mol C : 0.3386 mol H : 0.2033 mol O;

Divide each molar amount by the lesser of the three:

n(C) : n(H) : n(O) = 1.333 : 1.666 : 1 = 4 : 5 : 3


Since the resulting ratio is 4 C : 5 H : 3 O, the empirical formula is C4H5O3.


Answer: The empirical formula is C4H5O3.